# Extended Spin-Coherence Time in Strongly-Coupled Spin Baths in Quasi Two-Dimensional Layers

### v1

@@ I Detailed description of the system
In this section, we discuss the NV-center spin system, the interaction between the NV-center and the P1-center bath, describe details about the P1-center spin bath and, at the end, present the total Hamiltonian.
I.1 NV-center
We consider ensembles of NV-centers in [111]-oriented isotopically purified ${}^{12}$ diamond with a sparse disordered P1-center spin bath of concentration $\rho_{\text{P1}}$. Due to relatively long distances between the P1-centers, all defects can be treated as localized spins interacting via dipolar interaction. The diamond sample is placed in a strong magnetic field $B_{z}$ along the $z$-direction, which is also the [111]-direction, such that the three energy levels of the NV electron spin-1 are well separated (the NV-center nuclear spin is neglected, see below). In presence of this magnetic field $B_{z}$, the internal Hamiltonian of the NV electron spin in the laboratory reference frame reads ${\cal H}^{\text{lab}}_{\text{NV}}=DS_{z}^{2}-\gamma B_{z}S_{z}$, where $D$ denotes the zero-field splitting, $S_{z}$ is the z-projection spin-1 operator and $\gamma$ the electron-spin gyromagnetic ratio. The NV spin is initialized in $|\psi(0)\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|-1\rangle)$, where $|0\rangle$ and $|-1\rangle$ are eigenstates of $S_{z}$ with corresponding eigenvalues. The state $|+1\rangle$ remains unoccupied on the timescales considered because the $z$-component of the NV spin is conserved by ${\cal H}^{\text{lab}}_{\text{NV}}$ and the interaction Hamiltonian with the P1-center spin bath $\cal H_{\text{int}}$ (see below). The NV spin can thus be effectively treated as a spin-1/2 with the adapted $z$-projection operator
@@ I.1 NV-center
The ${}^{14}$ nuclear spin of the NV center can be assumed static on the timescales considered. The dominating interaction term in the strong magnetic field $B_{z}$ is of the form $I_{z}\tilde{S}_{z}$ [1], where $\tilde{S}_{z}$ is the $z$-projection operator of the nuclear spin. Such interaction would create an effective static magnetic field along the $z$-direction for the NV electron spin. Therefore, the effect of the NV nuclear spin on the NV electron spin is cancelled in the Hahn-echo sequence. Hence, we do not include the ${}^{14}$ nuclear spin in our numerical simulations.
@@ I.3 P1-center spin bath
A P1-center consists of an electron spin-1/2 and a ${}^{14}$ nuclear spin-1. Between the P1-centers, the interaction is dominated by the dipolar interaction between the respective electron spins. The corresponding Hamiltonian ${\cal H}_{\text{P1}}$ is given in Eq. (1) of the main text.
For each P1-center, the strong hyperfine interaction in the rotating reference frame is of the form
@@
where $A_{j}$ is the hyperfine interaction constant and $S_{jz}$ is the z-projection operator for the nuclear spin-1 of $j$-th P1-center [10, 48]. The hyperfine interaction constant $A_{j}$ depends on the orientation of the Jahn-Teller (JT) axis of the P1-center: for the [111]-axis (corresponding to the $z$-axis here), $A_{j}=114$ MHz, and $A_{j}=86$ MHz for the other three possible axis alignments. For the P1-center concentrations $\rho_{\text{P1}}$ considered here, the energy splittings induced by the hyperfine interaction are much larger than the dipole-interaction constants (see below) and, therefore, two P1-center electron spins can flip-flop only if the orientation of the corresponding nuclear spins and the JT axes is the same for the two P1-centers. Hence, the P1-center bath is effectively divided into five spin subgroups depending on the orientation of both, the JT axes and the nitrogen nuclear spins. Any P1-center remains in the same subgroup on the timescales of our simulations, because the nuclear spin dynamics as well as the JT reorientation have significantly longer characteristic timescales [60]. This complex structure of the P1-center bath implies that the majority of nearest neighbors of any P1-center spin is likely to belong to a different subgroup. These nearest neighbors create strong disordered local fields for the P1-center spin by means of the first Hamiltonian term in Eq. (1) but do not flip-flop with it.
@@ I.4 Total Hamiltonian
where ${\cal H}_{\text{NV}}$ describes the NV-center, $\cal H_{\text{int}}$ describes the interaction between the NV center and the P1-center bath, $\cal H_{\text{P1}}$ describes the interaction between the P1-centers and ${\cal H}_{\text{hf}}$ describes the hyperfine interaction for each P1-center. Substituting the individual terms in the expression above, the total Hamiltonian ${\cal H}_{\text{tot}}$ in the rotating reference frame reads
@@
For the dynamics in a Hahn-echo sequence, the effect of the first term ${\cal H}_{\text{NV}}$ is cancelled and, hence, this term can be left out.
@@
II Cluster-correlation expansion (CCE) method
[…]
~$\widetilde{\cal L}_{ni}\equiv\frac{\langle M_{x}(2\tau)\rangle_{ni}}{\prod_{{\cal B}\subset{\cal C}_{ni}}\widetilde{\cal L}_{\cal B}},$ (S8)
@@ II Cluster-correlation expansion (CCE) method
III Internal mean-field averaging
In this section, we discuss another type of mean-field bath-spin averaging often used for the suppression of unphysical CCE results (see, for example, Refs. [52, 53, 54]).
Internal mean-field averging is defined as follows: each term ~$\widetilde{\cal L}_{ni}$ is averaged over mean-field spin configurations separately before the individual factors are put together as shown in Eqs. (S8) and (S9). When using the internal averaging, correlations between the individual factors on the right-hand-side of Eqs. (S8) and (S9) are neglected. However, for disordered spin baths, these correlations seems to be negligible in practical calculations, cf. Fig. (2).
The internal averaging is particularly useful in CCE calculations with many unphysical terms, i.e., many clusters with $\langle M_{x}(2\tau)\rangle_{ni}>1$. When using internal averaging, the numerator and each term in the denominator in Eq. (S8) are averaged separately and this leads to the suppression of unphysical terms. The two kinds of averaging (normal and internal) can be combined in the following way. When obtaining the result for $\langle M_{x}(2\tau)\rangle$ in Eq. (S9) with internal averaging, this calculation can be repeated many times, thereby averaging over results for $\langle M_{x}(2\tau)\rangle$. We use both types of mean-field averaging in this letter as indicated below.
@@ III Internal mean-field averaging
IV Comparison of results of the pCCE method and exact simulations for small systems
To study the convergence behaviour of the pCCE method, we calculate the spin coherence $\langle M_{x}(2\tau)\rangle$ of NV spins in small baths of 20 electron spins-1/2. Here, we discuss results for three different spatial spin distributions, a two-dimensional spin lattice and two random three-dimensional baths, and compare the numerical pCCE results with results of an exact simulation. The exact simulations were performed using the Suzuki-Trotter method [56].
Figure S2: Spin coherence $\langle M_{x}(2\tau)\rangle$ decay of the NV-electron spin in three different small baths of 20 electron spins-1/2: (a) a two-dimensional spin lattice, and (b) - (c) three-dimensional spin baths with random spin distributions at concentration 1 ppm (see text). In all three cases, the pCCE results converge towards the exact results when increasing $K$. To demonstrate the effect of internal mean-field averaging, we also include results with and without internal mean-field average. Results obtained with internal averaging are labeled as “int. avg.”. In (b), already the CCE2 (equivalent to pCCE(2,1)) results show good agreement with exact results, whereas, in (c), there are sizable deviations between CCE2 and the exact solution.
For the exact simulations and also for the pCCE method with $K\approx 4$, we use canonical typicality for the P1-center bath state, which means that randomly chosen quantum states typically lead to good approximations of infinite-temperature expectation values [61, 62] for the bath spins. The calculations are averaged over sufficiently many randomly chosen initial quantum states for the P1-center bath to ensure convergence.
V Partitioning method
A key component of the proposed pCCE method is the partitioning of the spin bath into local strongly-interacting subsystems (partitions). We choose for simplicity the bath partitions to be of the same size $K$, corresponding to the same number of spins in the partitions. The partitioning is performed using a variant of the k-means algorithm. However, the method in general also allows a flexible partitioning with partitions of different size.
@@ V Partitioning method
The pCCE method is compatible with the gCCE approach (see, for example, Ref. [54]), which includes the central spin in each cluster. In fact, in order to use canonical typicality mentioned above, we used in our calculations the gCCE approach which is equivalent to the conventional CCE method for the systems considered.

### v2

@@ I Detailed description of the system
In this section, we discuss the NV-center spin system, the interaction between the NV center and the P1-center bath, describe details about the P1-center spin bath and, at the end, present the total Hamiltonian.
I.1 NV center
We consider ensembles of NV centers in [111]-oriented isotopically purified 12C diamond layers with a sparse disordered P1-center spin bath of concentration $\rho_{\text{P1}}$. Due to relatively long distances between the P1 centers, all defects can be treated as localized spins interacting via dipolar interaction. The diamond sample is placed in a strong magnetic field $B_{z}$ along the $z$-direction, which is also the [111]-direction, such that the three energy levels of the NV electron spin-1 are well separated (the NV-center nuclear spin is neglected, see below). It is important to note that the diamond layers are perpendicular to the [111]-direction and the magnetic field $B_{z}$. In presence of this magnetic field $B_{z}$, the internal Hamiltonian of the NV electron spin in the laboratory reference frame reads ${\cal H}^{\text{lab}}_{\text{NV}}=DS_{z}^{2}-\gamma B_{z}S_{z}$, where $D$ denotes the zero-field splitting, $S_{z}$ is the z-projection spin-1 operator and $\gamma$ the electron-spin gyromagnetic ratio. The NV spin is initialized in $|\psi(0)\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|-1\rangle)$, where $|0\rangle$ and $|-1\rangle$ are eigenstates of $S_{z}$ with corresponding eigenvalues. The state $|+1\rangle$ remains unoccupied on the timescales considered because the $z$-component of the NV spin is conserved by ${\cal H}^{\text{lab}}_{\text{NV}}$ and the interaction Hamiltonian with the P1-center spin bath $\cal H_{\text{int}}$ (see below). The NV spin can thus be effectively treated as a spin-1/2 with the adapted $z$-projection operator
@@ I.1 NV center
The 14N nuclear spin of the NV center can be assumed static on the timescales considered. The dominating interaction term in the strong magnetic field $B_{z}$ is of the form $I_{z}\tilde{S}_{z}$ [1], where $\tilde{S}_{z}$ is the $z$-projection operator of the nuclear spin. Such interaction would create an effective static magnetic field along the $z$-direction for the NV electron spin. Therefore, the effect of the NV nuclear spin on the NV electron spin is canceled in the Hahn-echo sequence. Hence, we do not include the 14N nuclear spin in our numerical simulations.
@@ I.3 P1-center spin bath
A P1 center consists of an electron spin-1/2 and a 14N nuclear spin-1. Between the P1 centers, the interaction is dominated by the dipolar interaction between the respective electron spins. The corresponding Hamiltonian ${\cal H}_{\text{P1}}$ is given in Eq. (1) of the main text.
For each P1 center, the strong hyperfine interaction in the rotating reference frame is of the form
@@
where $A_{j}$ is the hyperfine interaction constant and $S_{jz}$ is the z-projection operator for the nuclear spin-1 of $j$-th P1 center [10, 50]. The hyperfine interaction constant $A_{j}$ depends on the orientation of the Jahn-Teller (JT) axis of the P1 center: for the [111]-axis (corresponding to the $z$-axis here), $A_{j}=114$ MHz, and $A_{j}=86$ MHz for the other three possible axis alignments. For the P1-center concentrations $\rho_{\text{P1}}$ considered here, the energy splittings induced by the hyperfine interaction are much larger than the dipole-interaction constants (see below) and, therefore, two P1-center electron spins can flip-flop only if the orientation of the corresponding nuclear spins and the JT axes is the same for the two P1 centers. Hence, the P1-center bath is effectively divided into five spin subgroups depending on the orientation of both, the JT axes and the nitrogen nuclear spins. Any P1 center remains in the same subgroup on the timescales of our simulations, because the nuclear spin dynamics as well as the JT reorientation have significantly longer characteristic timescales [65]. This complex structure of the P1-center bath implies that the majority of nearest neighbors of any P1-center spin is likely to belong to a different subgroup. These nearest neighbors create strong disordered local fields for the P1-center spin by means of the first Hamiltonian term in Eq. (1) but do not flip-flop with it.
@@ I.4 Total Hamiltonian
where ${\cal H}_{\text{NV}}$ describes the NV center, $\cal H_{\text{int}}$ describes the interaction between the NV center and the P1-center bath, $\cal H_{\text{P1}}$ describes the interaction between the P1 centers and ${\cal H}_{\text{hf}}$ describes the hyperfine interaction for each P1 center. Substituting the individual terms in the expression above, the total Hamiltonian ${\cal H}_{\text{tot}}$ in the rotating reference frame reads
@@
For the dynamics in a Hahn-echo sequence, the effect of the first term ${\cal H}_{\text{NV}}$ is canceled and, hence, this term can be left out.
@@
II Mean-field averaging
@@
Another type of the mean-field bath-spin averaging is often used for the suppression of unphysical CCE results (see, for example, Refs. [54, 55, 56]), which we refer to as internal mean-field averaging in the following. Internal mean-field averaging is defined as follows: each term ~$\widetilde{\cal L}_{ni}$ (for definition, see main article) is averaged over mean-field spin configurations separately before the individual factors are put together in Eq. (S8). When using the internal averaging, correlations between the individual factors on the right-hand-side of Eq. (S8) are neglected. However, for disordered spin baths, these correlations seems to be negligible in practical calculations, cf. Fig. (2).
The internal averaging is particularly useful in CCE calculations with many unphysical terms, i.e., many clusters with $\langle M_{x}(2\tau)\rangle_{ni}>1$. When using internal averaging, the numerator and each term in the denominator in Eq. (2) are averaged separately and this leads to the suppression of unphysical terms. The two kinds of averaging (normal and internal) can be combined in the following way. When obtaining the result for $\langle M_{x}(2\tau)\rangle$ in Eq. (S8) with internal averaging, this calculation can be repeated many times, thereby averaging over results for $\langle M_{x}(2\tau)\rangle$. We use both types of mean-field averaging in this letter as explained below.
@@
II.1 Internal mean-field averaging in the conventional CCE method
We apply CCE3 to the bath of 20 electron spins-1/2 in a quasi-two-dimensional layer considered in Fig. 2 of the main article. We vary the number of samples for internal averaging to mitigate the unphysical behavior $\langle M_{x}(2\tau)\rangle>1$; the results are shown in Fig. S2. The results imply that, with increasing number of samples for the internal mean-field average, no enhancement of the accuracy is achieved because the results do not converge even after averaging over 1000 samples. We conclude that this small bath of pure electron spins-1/2 in quasi two dimensions represents a challenging system for the conventional CCE method requiring an even larger number of samples for the mean-field average.
Figure S2: Results of CCE3 applied to the 20 electron spin-1/2 bath in a quasi-two-dimensional layer (from Fig. (2)) for different numbers of samples for the internal average as indicated in the legend. At 1000 samples for the internal average, the CCE3 results have not converged and no signatures of convergence can be observed.
Number of spins Occurrence of $\langle M_{x}(2\tau)\rangle$ > 1
5 0 %
10 28 %
15 26 %
20 40 %
Table S1: Statistics for the occurrence of unphysical behavior ($\langle M_{x}(2\tau)\rangle>1$) for CCE3. We consider a series of smaller systems that we obtain by retaining only the 5, 10 or 15 closest spins to the NV center from the system of 20 electron spins in a quasi-two-dimensional layer, shown in Fig. 2 of the main article. These calculations were repeated 50 times and the fraction indicated in this table refers to this number of repetitions.
We assume that the problem of unphysical behavior ($\langle M_{x}(2\tau)\rangle>1$) becomes worse with increasing system size because, in larger spin baths, even more clusters are formed, each of which can lead to $\langle M_{x}(2\tau)\rangle>1$. To challenge this assumption, we consider a series of smaller systems that we obtain by retaining only the 5, 10 or 15 closest spins to the NV center from the system of 20 electron spins considered in Fig. S2. We use CCE3, fix the number of samples for the internal mean-field average to 250 and count the occurrence of unphysical behavior, i.e., $\langle M_{x}(2\tau)\rangle>1$, while repeating the same calculation with random choices of the mean-field bath-spin configurations. The probability for the occurrence of such behavior increases with larger system size as shown in Table S1.
III Partitioning method
A key component of the proposed pCCE method is the partitioning of the spin bath into local strongly interacting subsystems (partitions). We choose for simplicity the bath partitions to be of the same size $K$, corresponding to the same number of spins in the partitions. The partitioning is performed using a variant of the k-means algorithm. However, the method in general also allows a flexible partitioning with partitions of different size.
@@ III Partitioning method
The pCCE method is compatible with the gCCE approach (see, for example, Refs. [56, 38]), which includes the central spin in each cluster. In fact, in order to use canonical typicality mentioned above, we used in our calculations the gCCE approach, which is equivalent to the conventional CCE method for the systems considered.
@@
IV Structural comparison between pCCE and CCE
In this section, we compare the pCCE and CCE methods regarding the clusters included in the calculation. We start with an explicit example, which is followed by a more general discussion.
IV.1 An explicit example
[…]
For CCE1, we obtain
$\langle M_{x}(2\tau)\rangle_{\text{CCE1}}=\langle\langle\widetilde{\cal L}_{(1)}\widetilde{\cal L}_{(2)}\widetilde{\cal L}_{(3)}\widetilde{\cal L}_{(4)}\widetilde{\cal L}_{(5)}\widetilde{\cal L}_{(6)}\rangle\rangle,$ (S10)
where the contribution of each individual spin is included ~$\widetilde{\cal L}_{(n_{1})}=\langle M_{x}(2\tau)\rangle_{(n_{1})}$. In CCE2, the contribution of all possible spin pairs is also taken into account
$\begin{split}\langle M_{x}(2\tau)\rangle_{\text{CCE2}}=\langle\langle\widetilde{\cal L}_{(1)}\widetilde{\cal L}_{(2)}\widetilde{\cal L}_{(3)}\widetilde{\cal L}_{(4)}\widetilde{\cal L}_{(5)}\widetilde{\cal L}_{(6)}\widetilde{\cal L}_{(1,2)}\widetilde{\cal L}_{(1,3)}\widetilde{\cal L}_{(1,4)}\widetilde{\cal L}_{(1,5)}\widetilde{\cal L}_{(1,6)}\widetilde{\cal L}_{(2,3)}\widetilde{\cal L}_{(2,4)}\\ \widetilde{\cal L}_{(2,5)}\widetilde{\cal L}_{(2,6)}\widetilde{\cal L}_{(3,4)}\widetilde{\cal L}_{(3,5)}\widetilde{\cal L}_{(3,6)}\widetilde{\cal L}_{(4,5)}\widetilde{\cal L}_{(4,6)}\widetilde{\cal L}_{(5,6)}\rangle\rangle,\end{split}$ (S11)
where the two-spin terms read ~$\widetilde{\cal L}_{(n_{1},n_{2})}=\frac{\langle M_{x}(2\tau)\rangle_{(n_{1},n_{2})}}{\langle M_{x}(2\tau)\rangle_{(n_{1})}\langle M_{x}(2\tau)\rangle_{(n_{2})}}$, i.e., the two-spin terms must be normalized by the single-spin contributions.
In pCCE(1,2), only three individual partitions of size $K=2$ are included
$\langle M_{x}(2\tau)\rangle_{\text{pCCE(1,2)}}=\langle\langle\widetilde{\cal L}^{\prime}_{(1,2)}\widetilde{\cal L}^{\prime}_{(3,4)}\widetilde{\cal L}^{\prime}_{(5,6)}\rangle\rangle,$ (S12)
where the individual terms do not need to be normalized because the partitions are the basic units for pCCE, i.e., ~$\widetilde{\cal L}^{\prime}_{(n_{1},n_{2})}=\langle M_{x}(2\tau)\rangle_{(n_{1},n_{2})}$. This is why we can rewrite Eq. (S12) as
$\langle M_{x}(2\tau)\rangle_{\text{pCCE(1,2)}}=\langle\langle\widetilde{\cal L}_{(1)}\widetilde{\cal L}_{(2)}\widetilde{\cal L}_{(3)}\widetilde{\cal L}_{(4)}\widetilde{\cal L}_{(5)}\widetilde{\cal L}_{(6)}\widetilde{\cal L}_{(1,2)}\widetilde{\cal L}_{(3,4)}\widetilde{\cal L}_{(5,6)}\rangle\rangle,$ (S13)
which means that pCCE(1,2) goes beyond CCE1 (cf. Eq. (S10)) by including the most important two-spin terms.
In pCCE(2,2), the corresponding expansion reads
$\langle M_{x}(2\tau)\rangle_{\text{pCCE(2,2)}}=\langle\langle\widetilde{\cal L}^{\prime}_{(1,2)}\widetilde{\cal L}^{\prime}_{(3,4)}\widetilde{\cal L}^{\prime}_{(5,6)}\widetilde{\cal L}^{\prime}_{(1,2,3,4)}\widetilde{\cal L}^{\prime}_{(1,2,5,6)}\widetilde{\cal L}^{\prime}_{(3,4,5,6)}\rangle\rangle,$ (S14)
where
~$\widetilde{\cal L}^{\prime}_{(n_{1},n_{2},n_{3},n_{4})}=\frac{\langle M_{x}(2\tau)\rangle_{(n_{1},n_{2},n_{3},n_{4})}}{\langle M_{x}(2\tau)\rangle_{(n_{1},n_{2})}\langle M_{x}(2\tau)\rangle_{(n_{3},n_{4})}}.$ (S15)
In the denominator of the above expression, only the partitions forming the four-spin cluster $(n_{1},n_{2},n_{3},n_{4})$ enter, while, in the conventional CCE, a cluster of 4 spins needs to be normalized with all possible subclusters including clusters containing 1, 2 and 3 spins. An important implication of this modified normalization is that we obtain similar to Eq. (S13) now
$\begin{split}\langle M_{x}(2\tau)\rangle_{\text{pCCE(2,2)}}=\langle\langle\widetilde{\cal L}_{\text{CCE2}}\widetilde{\cal L}_{(1,2,3)}\widetilde{\cal L}_{(1,2,4)}\widetilde{\cal L}_{(1,3,4)}\widetilde{\cal L}_{(2,3,4)}\widetilde{\cal L}_{(3,4,5)}\widetilde{\cal L}_{(3,4,6)}\widetilde{\cal L}_{(3,5,6)}\widetilde{\cal L}_{(4,5,6)}\\ \widetilde{\cal L}_{(1,2,5)}\widetilde{\cal L}_{(1,2,6)}\widetilde{\cal L}_{(1,5,6)}\widetilde{\cal L}_{(2,5,6)}\widetilde{\cal L}_{(1,2,3,4)}\widetilde{\cal L}_{(3,4,5,6)}\widetilde{\cal L}_{(1,2,5,6)}\rangle\rangle,\end{split}$ (S16)
where ~$\widetilde{\cal L}_{\text{CCE2}}$ corresponds to the terms in Eq. (S11) entering the expression on the right-hand side within the brackets. Thus, by calculating the contribution of a four-spin cluster, also the contribution of the three-spin clusters are included without their explicit calculation.
[…]
IV.2 Comparison of the number of clusters
The total time $T_{\text{tot}}$ required to obtain the final results for CCE or pCCE can be estimated as follows
$T_{\text{tot}}\approx N_{\text{sys}}\cdot N_{\text{clusters}}\cdot N_{\text{average}}\cdot T(\text{size of largest clusters}),$ (S17)
where $N_{\text{sys}}$ is the number of random spin distributions considered, $N_{\text{clusters}}$ is the number of largest clusters of the expansion, $N_{\text{average}}$ is the number of samples for the mean-field average and $T(\text{size of largest clusters})$ is the time required to calculate the contribution for one of the largest clusters. Only the largest clusters are considered in the above estimate because the calculation of their contribution usually covers most of the overall calculation time. While the factor $N_{\text{sys}}$ is the same for both methods, pCCE offers advantages in the remaining three factors entering Eq. (S17) as we show in the following.
To calculate the number of clusters $N_{\text{clusters}}$, we make the assumption that the dipole radius $r_{\text{d}}$ (defined below) is infinite, i.e., any spins can be combined in a cluster irrespective of the distance between the spins. Then, the number of largest clusters of size $NK$ is
$\begin{split}N_{\text{clusters}}=&\binom{\frac{N_{s}}{K}}{N}\text{ for pCCE($N$,$K$)}\\ N_{\text{clusters}}=&\binom{N_{s}}{NK}\text{ for CCE$NK$},\end{split}$ (S18)
where $N_{s}$ is the total number of bath spins. For the typical case, when $N_{s}\gg KN$, $\binom{N_{s}/K}{N}\ll\binom{N_{s}}{NK}$. For $N_{s}=100$, we provide the numbers of clusters of different sizes included in pCCE and CCE in Table S2. For the pCCE method, we also indicate how many clusters are effectively included without calculating their contribution separately, as explained in the previous section. For example, using pCCE(2,4), in addition to the 300 clusters of size 8, a large number of smaller clusters is included, cf. last column of Table S2.
Cluster size CCE4 CCE8 pCCE(2,2) pCCE(2,2) pCCE(2,4) pCCE(2,4)
calculated calculated calculated included calculated included
1 100 100 - 100 - 100
2 4 950 4 950 50 4950 - 4950
3 161 700 161 700 - 4900 - 14500
4 3 921 225 3 921 225 1225 1225 25 20425
5 - $7.53\cdot 10^{7}$ 16800
6 - $1.19\cdot 10^{9}$ 8400
7 - $1.60\cdot 10^{10}$ 2400
8 - $1.86\cdot 10^{11}$ 300 300
Table S2: Number of clusters of different size (number of spins included) calculated in CCE and pCCE for a bath of $N_{s}=100$ spins. For CCE, the number of calculated and incorporated clusters is the same (see text). In pCCE(2,$K$), only clusters of size $K$ and $2K$ are calculated but many other clusters are effectively included.
More importantly, when comparing the number of four-spin clusters for CCE4 and the number of eight-spin clusters for pCCE(2,4), the overall calculation time is expected to be smaller for pCCE(2,4) than for CCE4 because the ratio of these numbers is approximately $10^{4}$, cf. Table S2, which is much larger than the expected scaling of the computational time $4^{n}$ with the spin number $n$ ($2564^{4}=$ 256). This discrepancy in the number of clusters becomes even stronger, when $N$ is further increased. Even after introducing a finite dipole radius $r_{\text{d}}$ (see below) and, thereby, reducing the number of four-spin clusters for CCE4, we anticipate that the pCCE calculation terminates faster. We expect pCCE(2,4) to yield results comparable to CCE4 (or better) in a shorter time.
When considering $N_{\text{average}}$, pCCE also offers an advantage here, as shown above. For example, when comparing CCE4 and pCCE(2,4), we expect that CCE4 requires at least 100 times more samples for the mean-field average than pCCE(2,4) and this factor further grows when increasing $N$.
[…]
V Comparison of results of the pCCE method and exact simulations for small systems
To study the convergence behavior of the pCCE method, we calculate the spin coherence $\langle M_{x}(2\tau)\rangle$ of NV spins in small baths of 20 electron spins-1/2. Here, we discuss results for three different spatial spin distributions, a two-dimensional spin lattice and two random three-dimensional baths, and compare the numerical pCCE results with results of an exact simulation. The exact simulations were performed using the Suzuki-Trotter method [58].
Figure S4: Spin coherence $\langle M_{x}(2\tau)\rangle$ decay of the NV electron spin in three different small baths of 20 electron spins-1/2: (a) a two-dimensional spin lattice, and (b) - (c) three-dimensional spin baths with random spin distributions at concentration 1 ppm (see text). In all three cases, the pCCE results converge towards the exact results when increasing $K$. To demonstrate the effect of internal mean-field averaging, we also include results with and without internal mean-field average. Results obtained with internal averaging are labeled as “int. avg.”. In (b), already the CCE2 (equivalent to pCCE(2,1)) results show good agreement with exact results, whereas, in (c), there are sizable deviations between CCE2 and the exact solution.
For the exact simulations and also for the pCCE method with $K\approx 4$, we use canonical typicality for the P1-center bath state, which means that randomly chosen quantum states typically lead to good approximations of infinite-temperature expectation values [59, 60, 61] for the bath spins. The calculations are averaged over sufficiently many randomly chosen initial quantum states for the P1-center bath to ensure convergence.